OpenStudy (anonymous):

Write in polar form -1+i√3

OpenStudy (anonymous):

\(\sqrt{(-1)^2+(\sqrt{3})^2}(\cos(\arctan(-\sqrt3)+i \sin(\arctan(-\sqrt3))\)

OpenStudy (anonymous):

@myko YOU ARE GENUIS

OpenStudy (anonymous):

lol